OpenStudy (anonymous):

Write the equation in slope-intercept form. What are the slope and y-intercept? -11x + 9y = -12 A. y = 11/9x - 4/3 slope: 11/9 y-intercept: 4/3 B. y = -11/9x + 4/3 slope: 11/9 y-intercept: 4/3 C. y = 11/9x - 4/3 slope: 11/9 y-intercept: -4/3 D. y = 11/9x + 4/3 slope: 4/3 y-intercept: 11/9

OpenStudy (anonymous):

First step: rewriting -11x + 9y = -12 9y = 11x - 12 (getting the "x" term over to the right) y = (11/9)x - 4/3 (diving both sides by 9) Second step: slope-intercept form is: y = mx + b so, matching up, we have: (11/9) for slope and -4/3 for the intercept Answer is "c"

OpenStudy (anonymous):

Is this making sense to you now? @HannaaLovee

OpenStudy (anonymous):

Good luck in all of your studies and thx for the recognition! @HannaaLovee

OpenStudy (anonymous):

thank you very much ! :)

OpenStudy (anonymous):

You are entirely welcome! Nice working with you!